Calculate the magnitude of the electric field at one corner of a square 1.03 m ona side if the other three corners are occupied by 5.03x10-6 C charges. - 21307.21 N/C
Answer #1

field due to either adjacent charge

E1 = E2 = k q /r^2

E1 = E2 = 9*10^9* 5.03*10^-6 / 1.03^2

E1 = E2 = 42671.32 N/C

field due to corner charge

E3 = 9*10^9* 5.03*10^-6 / ( 1.03* sqrt(2))^2

E3 = 21335.66 N/C

net horizontal field

Ex = E1 + E3 cos 45 = 57758 N/C

net vertical field

Ey = E2 + E3 sin 45 = 57758 N/C

net field

E^2 = Ex^2 + Ey^2

E = 81682 N/C

=========

do comment in case any doubt, will reply for sure.. Goodluck